Suppose V(r) = 4/3(pi)r^3 and p(t)= V(3t). Evaluate and simplify p(3)
Answer by stanbon(75887)

Suppose V(r) = 4/3(pi)r^3 and p(t)= V(3t). Evaluate and simplify p(3)
------
P(3) = V(3*3) = V(9) = (4/3)pi*9^3 = (4*9^2*3)pi = 972pi
===============================================
Cheers,
Stan H.